OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 3x2 + 5x + 1 = 0 I got x=-5 (plus or minus) the sqrt of 13/2 and apparently my answer is wrong

OpenStudy (anonymous):

\(a=3,b=5,c=1\)

OpenStudy (anonymous):

yes, I had that as a b and c

OpenStudy (anonymous):

\[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\frac{-5\pm\sqrt{5^2-4\times 3\times 1}}{2\times 3}\]

OpenStudy (anonymous):

i see the mistake denominator should be 6 and not 2

OpenStudy (anonymous):

okay thank you! @satellite73

OpenStudy (anonymous):

yw